OpenStudy (anonymous):

Essay. Show your work. Find the following product: (x^2-5)(x^3+4x+7)

jhonyy9 (jhonyy9):

- hope that you will can makeing the calcules - so you need just multiplie term by terms from the first parantheses with the terms from second parantheses ,ok ?

jhonyy9 (jhonyy9):

- first step x^2 * x^3 = ?

jhonyy9 (jhonyy9):

- are you there ?

jhonyy9 (jhonyy9):

courage ,i will help you

OpenStudy (anonymous):

\[\begin{align} &&\large x^3+4x+7\\ &\times &\large x^2-5\\ \hline \\ &&\large -5x^3-20x-35\\ &+&\large x^5+4x^3+7x^2\\ \hline \\ &&\large x^5-x^3+7x^2-20x-35 \end{align}\]